Brackish water is desalinated by evaporation for irrigation purposes. If the feed water has 800.96 ppm of salt, and if the product has to contain 0.0034 % salts as illustrated in Fig 1, what fraction of the feed water should bypass the evaporator?


1
Expert's answer
2021-10-07T03:14:41-0400



feed water = 800.96 ppm

product water = 0.0034% = 34 ppm


"800.96 ×( 1-x)= 34 × x\\
800.96- 800.96x = 34x\\
800.96 = 834.96x\\
x = 800.96/834.96 = 0.959"


part needed to bypass = "1-0.959" =

"0.041"